Find four consecutive integers whose sum is -38.
Found 2 solutions by mananth, Alan3354:
Answer by mananth(16949)

Find four consecutive integers whose sum is -38.
Let the numbers be x, x-1, x-2, x-3
The sum of all these is -38
4x-6=-38
4x=-38+6
4x=-32
X=-8
The numbers are -8 , -9 -10, -11
M. Ananth

Answer by Alan3354(69443)

The easy way:
-38/4 = -9.5 which is the center.
2 numbers are adjacent on either side.
--> -11, -10, -9, -8